OpenStudy (anonymous):

One of the roots of the quadratic equation x^2-2kx-32k=0, where k is a constant and k is not=0,is one fourth the other root.Find the possible value of k.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Say \(\rm \alpha , ~ \frac{\alpha }{4}\) are the roots of given eqn

OpenStudy (anonymous):